A3. A direct tensile stress σ,-160 MN/m2 and a direct compressive stress σ,--120 MN/m2 are maximum principal stress in the material is limited to 200 MN/m2 (in tension). Use Mohrs circle to find the allowable value of shear stress at the point. [10 Marks] Final Answer: Max shear stress-radius of circle-180 MN/m
Answer #1

We know that the principal stress is given by
の
On putting the value
G) 1(160-120) + V/(160-(-120))2+4制 200 =
On solving we get
Try- 113.14 MN/m2
Now drawing the point A and B on the mohr circle
Diameter of mohr circle
уп60_ (-120)]2 + [113-(-113)12
= 360
Now the maximum shear stress = Radius of mohr circle = 360/2 = 180 MN/m2
(160, 113) (-120,-113)